(i) Show that the sequence {n/(2n-1)}_(n=1)^∞ is monotonic. (ii) Is it bounded above and/or below? Give a reason for your answer. (iii) Can you concluded that it converges? Give a reason for your answer.


1
Expert's answer
2021-05-10T15:40:35-0400

"i)\newline

Given, a_n=\frac{n}{2n-1}.\newline
a_n-a_{n+1}=\frac{n}{2n-1}-\frac{n+1}{2n+1}.\newline
=\frac{2n^2-n-2n^2+n-2n+1}{(2n-1)(2n+1)}\newline
=\frac{-1}{2n+1}<0,\space \text{for all n}\newline
\implies a_n<a_{n+1}\newline
\implies a_n \text{is monotonic.}\newline

ii)\newline
a_1=\frac{1}{3}\newline
a_2=\frac{2}{5}\newline
a_3=\frac{3}{7}\newline
...\newline
Since, a_n is increasing sequence.\newline
Then, lim a_n=\frac{1}{2}.\newline
\text{Therefore, the given sequence islower bounded by }\frac{1}{3}\newline
\text{and upper bounded by 0.5.}\newline
iii)\newline
\text{Since, the given sequence is monotonic and bounded.}\newline
\text{Therefore, it is convergent.}"